Esquire Company needs to acquire a molding machine to be used in its manufacturing process. Two types of machines that would be appropriate are presently on the market. The company has determined the following: (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) Machine A could be purchased for $15,000. It will last 10 years with annual maintenance costs of $500 per year. After 10 years the machine can be sold for $1,575. Machine B could be purchased for $12,500. It also will last 10 years and will require maintenance costs of $2,000 in year three, $2,500 in year six, and $3,000 in year eight. After 10 years, the machine will have no salvage value. Required: Assume an interest rate of 8% properly reflects the time value of money in this situation and that maintenance costs are paid at the end of each year. Ignore income tax considerations. Calculate the present value of Machine A & Machine B. Which machine Esquire should purchase? (Negative amounts should be indicated by a minus sign. Do not round intermediate calculations. Round your final answers to nearest whole dollar amount.)
Answer #1
Question 1 :- Present value computation
Answer:- Refer below for the present value computation
Question 2 :- Which machine Esquire should purchase
Answer:- Based on the below present value computation the Cesquire should buy Machine B as it has lowest present value of cash flow
Year Nature of cash flow Cash out flow for Machine A Cash out flow for Machine B Present value factor @ 8% Present value of cash out flow for Machine A Present value of cash out flow for Machine B
A B C D E=1/(1+8%)^A F=C*E G=D*E
0 Cost of Machine             (15,000)        (12,500)             1.0000             (15,000.00)         (12,500.00)
1 Annual maintenance cost                  (500)             0.9259                 (462.96)                      -  
2 Annual maintenance cost                  (500)             0.8573                 (428.67)                      -  
3 Annual maintenance cost                  (500)          (2,000)             0.7938                 (396.92)           (1,587.66)
4 Annual maintenance cost                  (500)             0.7350                 (367.51)                      -  
5 Annual maintenance cost                  (500)             0.6806                 (340.29)                      -  
6 Annual maintenance cost                  (500)          (2,500)             0.6302                 (315.08)           (1,575.42)
7 Annual maintenance cost                  (500)             0.5835                 (291.75)                      -  
8 Annual maintenance cost                  (500)          (3,000)             0.5403                 (270.13)           (1,620.81)
9 Annual maintenance cost                  (500)             0.5002                 (250.12)                      -  
10 Annual maintenance cost                  (500)             0.4632                 (231.60)                      -  
10 Salvage value                 1,575             0.4632                   729.53                      -  
Present value of cash out flow                (17,626)            (17,284)